# An MHD Simulation of the Possible Modulations of Stellar CMEs Radio Observations by an Exoplanetary Magnetosphere

### v1

@@ Abstract
Type II radio bursts are the indicator of adverse space weather in a stellar system. These radio bursts are the consequence of shock wave acceleration due to the coronal mass ejection (CME). Here, we perform a series of magnetohydrodynamic (MHD) simulations of a CME-driven star-planet system in order to investigate the modulation in radio burst mechanism by a close-in exoplanetary system. We use a model for the stellar wind with a close-in exoplanet, and a CME model based on the eruption of a flux rope. We are able to generate synthetic radio burst images from our MHD simulations. We find that radio burst like phenomena is most likely to be observed for moderately active solar like stars and close-in exoplanetary systems have significant influence on the nature of radio burst spectrum. We find that when the planetary field is not too strong, the planetary magnetosphere is pushing against the CME, increasing its density so the radio burst is visible at higher frequencies. When the planetary field is very strong, the large magnetosphere does not leave room for the CME shock to evolve so the radio burst is more visible in the lower frequencies associated with the weak compression at the flanks of the CME shock. In case of highly active solar-like stars, strong overlying stellar fields weakens the solar-like CME shock, thus generates very weak (almost non-visible) radio burst signals. For HD 189733 (moderate stellar field), only intensity difference is visible when the CME arrives the planet. We also do not find significant modulation in the radio emission by a close-in exoplanet system when the stellar magnetic field is complex. In summary, our result suggests that the nature of the radio burst spectrum is highly dependent on the topology of the stellar magnetic field and the close-in exoplanetary magnetic field strength.
@@ 3.1 Star with Simple Dipolar Magnetic Field
3.2 Star with Very Strong Dipolar Magnetic Field: Is it possible to observe the radio burst for highly active stars?
@@
Our results already indicate the possibility of observing stellar radio bursts for moderately active stars. However, the question remains, is it possible to observe radio bursts for highly active stars? To investigate this issue, we consider a star with a very strong 100G dipolar magnetic field. We formulate two specific cases for this investigation - in the first scenario, we assume that there is no planet in the stellar system, and in the second scenario, we assume the presence of a planet near the central star. In the second scenario, we place a hot Jupiter-like planet at a 10 stellar radii distance apart from the central star.
@@
Figure 5 shows the dynamic radio spectrums obtained from our simulations of highly active stars. The top panel shows the dynamic radio spectrum when there is no planet near the star and the bottom panel shows the same when there is a planet near the star. In both scenarios, we do not find any strong signature of a radio burst. A radio burst is basically a sudden increase in the radio emission. Radio emission fluxes are not very high in both scenarios. Our results suggest that radio burst is very unlikely to be observed for highly active stars. Figure 5 shows the velocity evolution after one hours of the CME eruption. Left panel corresponds to the situation when there is no planet near the central star, while right panel corresponds to the situation when there is a planet near the central star. Comparison between two panels clearly shows that propagation of shock waves are impacted by the close-in exoplanets. However, second panel also indicates that CMEs are suppressed by strong overlying fields. Basically, CME related shocks are not able to propagate outwards quickly due to the strong overlying magnetic fields and become weaker very quickly.
@@
Type II radio bursts are the evidence for the presence of shocks in the stellar atmosphere. These shocks generally associated with the coronal mass ejection like events and can emit radio signals from submeter to kilometer wavelengths. Stronger shocks results in higher radio frequency generation in the stellar system. Our present simulation for highly active stars indicates the confinement of CME associated shocks by large scale overlying magnetic fields present in the stellar system. Shocks are weakened quickly and resulting in a very weak radio signal. Previous studies also indicate the suppression of CME like events by large scale magnetic fields and consequent brightening of solar corona in X-ray (see (Alvarado-Gómez et al. 2019)). Our result suggest that confinement of CME like events by strong overlying magnetic fields will lead to non-observance of radio burst like events. However, we note that our choice of CME energy is solar like (around $10^{33}$-$10^{34}$ ergs). It may be possible that CME energy is very high in case of highly active stars. In that scenario, stellar super CMEs may break the confinement and generate strong radio signals. However, we do not have any observation of super CMEs till now.
@@ 5 Conclusion
Observing and characterizing stellar coronal mass ejections and their interactions with the exoplanetary atmosphere are important for the understanding of stellar systems and their space weather. In this study, we perform CME-driven star-planet interaction MHD simulations to investigate the nature of the radio burst spectrum for stellar systems. Without direct observation of stellar CMEs, dynamic radio burst spectrums are believed to be an indirect proxy of stellar CMEs. In our study, we choose different surface magnetic field topologies for our central star. We sometimes place one hot Jupiter planet near our central star to investigate the interaction of stellar CMEs with the close-in exoplanets. In summary, we use a set of simulations aiming to demonstrate the feasibility of observing stellar coronal mass ejections and their interactions with the close-in exoplanets in radio wavelength.
@@
In the case of our simulations with highly active stars (the stellar surface magnetic field is very high), we do not find any clear signature of a radio burst. CME in our model setup is completely suppressed by strong-overlying stellar magnetic fields, making radio signals much weaker compared to the solar scenario. However, we note that the energy of the CMEs in our setup is solar-like. It is possible that a super-CME (very highly energetic) like events can break the strong overlying stellar magnetic fields and generate strong visible radio signals.
Additionally, we note that our study is based on the comparatively simpler stellar surface magnetic field topology. Highly complex coronal magnetic field structures (containing many coronal hole/open magnetic field regions) may have an impact on the nature of the radio burst spectrum. Our model also considers only thermal radio emission, not coherent (nonthermal) radio emission. Simultaneous observations in optical, radio, EUV, and X-ray are necessary for the detection and characterization of stellar CMEs. MHD simulations of CME drive star-planet interaction may help to constrain and understand multiwavelength observations of stellar CMEs.

### v2

@@ Abstract
Type II radio bursts are the indicator of adverse space weather in a stellar system. These radio bursts are the consequence of shock wave acceleration due to the coronal mass ejection (CME). In this study, we conduct a series of magnetohydrodynamic (MHD) simulations of CME-driven star–planet systems to investigate how close-in exoplanets modulate radio burst characteristics. We use a model for the stellar wind with a close-in exoplanet, and a CME model based on the eruption of a flux rope. We are able to generate synthetic radio burst images from our MHD simulations. We find that radio burst like phenomena is most likely to be observed for moderately active solar like stars and close-in exoplanetary systems have significant influence on the nature of radio burst spectrum. We find that when the exoplanet’s magnetic field is relatively weak, its magnetosphere compresses the CME plasma, increasing local density and shifting the radio emission to higher frequencies. Conversely, a strong planetary magnetic field results in a large magnetosphere that prevents effective CME- shock development, producing weaker radio emission concentrated at lower frequencies, particularly at the flanks of the CME. For highly active solar-like stars, strong overlying stellar magnetic fields suppress the CME shock, greatly diminishing radio burst visibility. For HD 189733 (moderate stellar field), only intensity difference is visible when the CME arrives the planet. We also do not find significant modulation in the radio emission by a close-in exoplanet system when the stellar magnetic field is complex. In summary, our findings highlight that the nature of the radio burst spectrum is strongly dependent on both the topology of the stellar magnetic field and the magnetic strength of close-in exoplanets.
@@ 3.1 Star with Simple Dipolar Magnetic Field
3.2 Star with Very Strong Dipolar Magnetic Field: Is it possible to observe the radio burst for highly active solar-like stars?
@@
Our results already indicate the possibility of observing stellar radio bursts for moderately active stars. However, the question remains, is it possible to observe radio bursts for highly active solar-like stars? To investigate this issue, we consider a solar-like star with a very strong 100G dipolar magnetic field. We formulate two specific cases for this investigation - in the first scenario, we assume that there is no planet in the stellar system, and in the second scenario, we assume the presence of a planet near the central star. In the second scenario, we place a hot Jupiter-like planet at a 10 stellar radii distance apart from the central star.
@@
Figure 5 shows the dynamic radio spectrums obtained from our simulations of highly active solar-like stars. The top panel shows the dynamic radio spectrum when there is no planet near the star and the bottom panel shows the same when there is a planet near the star. In both scenarios, we do not find any strong signature of a radio burst. A radio burst is basically a sudden increase in the radio emission. Radio emission fluxes are not very high in both scenarios. Our results suggest that radio burst is very unlikely to be observed for highly active solar-like stars. Figure 5 shows the velocity evolution after one hours of the CME eruption. Left panel corresponds to the situation when there is no planet near the central star, while right panel corresponds to the situation when there is a planet near the central star. Comparison between two panels clearly shows that propagation of shock waves are impacted by the close-in exoplanets. However, second panel also indicates that CMEs are suppressed by strong overlying fields. Basically, CME related shocks are not able to propagate outwards quickly due to the strong overlying magnetic fields and become weaker very quickly.
@@
Type II radio bursts are the evidence for the presence of shocks in the stellar atmosphere. These shocks generally associated with the coronal mass ejection like events and can emit radio signals from submeter to kilometer wavelengths. Stronger shocks results in higher radio frequency generation in the stellar system. Our present simulation for highly active stars indicates the confinement of CME associated shocks by large scale overlying magnetic fields present in the stellar system. Shocks are weakened quickly and resulting in a very weak radio signal. Previous studies also indicate the suppression of CME like events by large scale magnetic fields and consequent brightening of solar corona in X-ray (see (Alvarado-Gómez et al. 2019)). Our result suggest that confinement of CME like events by strong overlying magnetic fields will lead to non-observance of radio burst like events. However, we note that our chosen CME energy range is solar-like, around $10^{33}$-$10^{34}$ ergs. It is possible that in the case of highly active stars, CME energies could be significantly higher. In such scenarios, these stellar super-CMEs might overcome magnetic confinement and generate strong radio emissions. However, there is currently no observational evidence for the existence of such super-CMEs.
@@ 5 Conclusion
Observing and characterizing stellar coronal mass ejections and their interactions with the exoplanetary atmosphere is crucial for the understanding of stellar systems and their space weather. In this study, we perform magnetohydrodynamic (MHD) simulations of CME-driven star-planet interactions to investigate the nature of the resulting radio burst spectrum for stellar systems. Without direct observation of stellar CMEs, dynamic radio burst spectrums often considered as an indirect proxy of stellar CMEs. In our study, we choose different surface magnetic field topologies for our central star. We sometimes place one hot Jupiter planet near our central star to investigate the interaction of stellar CMEs with the close-in exoplanets. Our simulations aim to assess the feasibility of detecting stellar CMEs and their planetary interactions via radio wavelength observations.
@@
In the case of our simulations with highly active solar-like stars (the stellar surface magnetic field is very high), we do not find any clear signature of a radio burst. CME in our model setup is completely suppressed by strong-overlying stellar magnetic fields, making radio signals much weaker compared to the solar scenario. However, we note that the energy of the CMEs in our setup is solar-like. It is possible that super-CME events, with much higher energies, could overcome this magnetic suppression and produce detectable radio signatures–though such events have yet to be observed.
Additionally, we note that our study is based on the comparatively simpler stellar surface magnetic field topology. Highly complex coronal magnetic field structures (containing many coronal hole/open magnetic field regions) may have an impact on the nature of the radio burst spectrum. Our model also considers only thermal radio emission, not coherent (nonthermal) radio emission. Coordinated multiwavelength observations–spanning optical, radio, EUV, and X-ray–are essential for the detection and characterization of stellar CMEs. MHD simulations, such as those presented here, can provide valuable constraints and interpretation for such observations.